OpenStudy (anonymous):

What is the value of the expression 7^–3?

OpenStudy (mathmale):

Hello, Kiki! First, what does 7^(-3) mean to you?\[7^3~?\]

OpenStudy (anonymous):

\[7^{-3}\] as in negative 3.

OpenStudy (mathmale):

What are the values of the following?\[7^{-1},~7^{-2},~7^{-3}?\]

OpenStudy (anonymous):

I don't know, that's why I asked the question,

OpenStudy (mathmale):

Let's start with examples.\[2^1=2,~2^{2}4. ~~2^3=?\]

OpenStudy (anonymous):

8.

OpenStudy (mathmale):

2^3 = 8, yes. What about 2^0? What about 2^(-1)? How do you interpret a negative exponent?

OpenStudy (anonymous):

The first one would be 1. I think the second is 0.5. You interpret a negative exponent by dividing, doing the opposite of multiplying.

OpenStudy (mathmale):

2^0 is indeed 1. Very good.

OpenStudy (mathmale):